# Decode Neural signal as Speech

### v2

@@ 4.1 Dataset
Decoding Results on Schoffelen (Schoffelen et al., 2019)
@@
4.2 Implementation Details
Model modification. We adopt whisper base model (Radford et al., 2023) as our base model. To adapt the input size of the neural signal, we use two convolution1d layers, the kernel sizes are 3, and stride is 1,2 respectively to replace the first convolution layer in the whisper-base model of dimension 512.
Training. To implement eeg-to-text (Wang & Ji, 2022), we use the original word timestamps provided by GWilliams (Gwilliams et al., 2023) to segment MEG waves for each word feature. The training settings are the same as the origin code.
All models are trained on Nvidia H800 GPUs. We use a learning rate of 1e-3 and batch size of 64 for training 500 epochs and select the best model on evaluation loss. We use the AdamW as the optimizer for training all the models.
Evaluation. For non-teacher forcing generation configuration, we use the beam search with a beam size of 5 and set the repetition penalty to 5, with no repeat n-gram size to 2, which is very important to avoid repeating. In addition, we implement post-processing composed of filtering and lowering alphabets to improve performance.
4.3 Metrics
As shown in Table 1, we compare our results to input Gaussian noise to the NeuSpeech model and randomly select sentences from the test set and whisper base model speech transcription to give the understanding of up and down limits. We also compare previous method eeg-to-text (Wang & Ji, 2022).
Based on the information provided in Tab. 1, it is evident that our method outperforms randomly selected sentences, noise input, and eeg-to-text (Wang & Ji, 2022). NeuSpeech has demonstrated the feasibility of decoding text from non-invasive brain recordings. Prior to NeuSpeech, many previous works (Duan et al., 2023; Wang & Ji, 2022) in this field relied on teacher forcing to evaluate their models, Defossez et al. (Défossez et al., 2023) used classification accuracy. However, NeuSpeech introduces a more fair and real-world setting by employing end-to-end training, which directly ties neural signals to text. In comparison to previous works evaluated using teacher forcing, NeuSpeech has shown remarkable improvement, achieving BLEU-1, BLEU-2, BLEU-3, and BLEU-4 scores of 60.3, 55.26, 51.24, and 47.78, respectively, which is close to the performance of speech recognition.
To ensure a fair comparison, we also implement teacher forcing for both the eeg-to-text model (Wang & Ji, 2022) and NeuSpeech. It is evident that our BLEU-1 score of 80.20 is almost four times that of the eeg-to-text (Wang & Ji, 2022) model, highlighting the superiority of our approach. Furthermore, we have surpassed speech recognition methods with teacher forcing, emphasizing that our model can significantly improve performance by eliminating accumulated errors. It is worth noting that, at this stage, BLEU is the most suitable metric for evaluating the MEG-to-text system, as ROUGE and WER tend to yield high scores for noise and random models.
4.4 Generated Samples
Table 3: Transcription results comparison with different methods. Bold font means exact match of words. Underline for similar but not exact matching. w/ means with. tf means teacher-forcing, pt means pretraining, jt means joint-training. eeg-to-text (Wang & Ji, 2022) here is using beam search as the same setting of ours which is described in Sec. 4.2.
(1) GT Arthur replaced the tray onto the lacquered cherry table that stood in the center of his ordered chamber
NeuSpeech The rental car records pointed to lacquered cherry-schedung and stood in the center of his ordered chamber in motel from Acres comb in?
NeuSpeech w/ tf Arthur replaced thethe cartray onto thethe lacquered cherry-table inthat stood in thethe center of his ordered chamber in
NeuSpeech w/ pt Arthur replaced the tray onto the lacquered cherry table to match from Acres.
NeuSpeech w/ jt The rental car records pointed to Acres and stay in the center of his ordered chamber
eeg-to-text (Wang & Ji, 2022) He said he had no idea how long it would take him to get there
eeg-to-text (Wang & Ji, 2022) w/ tf was the old with the tablekeyered table table and had in the middle of the room table
@@ 4.4 Generated Samples
eeg-to-text (Wang & Ji, 2022) He said he had no idea how long it would take him to get back home
eeg-to-text (Wang & Ji, 2022) w/ tf father was running gone I was ready to the dark drive up Iacia me the window said me to was a a he was looking for
@@
eeg-to-text (Wang & Ji, 2022) It was the first time I had seen him in a long time and it made me very nervous
eeg-to-text (Wang & Ji, 2022) w/ tf little then he the seat and the place go it take it he one man would you
In Tab. 2, we present visual examples of text that have been generated from MEG signals. The training of Schoffelen (Schoffelen et al., 2019) is described in sec. 4.5.1 These examples showcase the capabilities of our model in generating meaningful text, despite the difficulties associated with translating thoughts and the limited prior research in this area. While our model may not yet achieve the same level of performance as nowadays speech recognition tasks, it demonstrates the ability to align keywords and construct sentences with similar structures. This is a promising indication of the potential of our approach in the field of brain-to-text translation.
@@
We have provided Tab. 3 below, which compares the results of different approaches, namely NeuSpeech, NeuSpeech with teacher forcing, NeuSpeech with pre-training, and joint training. We observe that using NeuSpeech alone often result in semantically coherent sentences, even if not all the words were correct. Particularly in example (3), there was even a hint of expressive tone. However, when NeuSpeech was generating text with teacher forcing, although there were more correct word matches, there were often consecutive word repetitions, such as ”thethe.” The reason for this occurrence is not yet clear, but it significantly affects the semantic coherence of the sentences.
@@
In contrast, the joint training model seems to tolerate longer sequences of word errors. For example, in (1), the model does not predict a single correct word for most of the sentence but accurately predicts the last seven words, which is neglected by the higher-scoring pre-trained model. This indicates that, even for a model that reaches a certain performance level, we cannot solely rely on scoring metrics. Each model may have unique attributes that we can exploit. In addition, previous approaches for EEG-to-text translation seems to be less effective, as even with the use of teacher forcing, they produced only a few words that match the true labels. This further highlights the superiority of our approach.
In summary, the analysis of generated samples from different approaches reveals valuable insights. NeuSpeech demonstrates the ability to generate semantically coherent sentences, while NeuSpeech with teacher forcing and pre-training offers improvements in word accuracy, albeit with some drawbacks. The joint training model shows resilience to long sequences of word errors. These findings emphasize the importance of considering both quantitative metrics and understanding the unique properties and strengths of different models. Furthermore, the comparison with previous EEG-to-text (Wang & Ji, 2022) methods underscores the advancements achieved by our approach.
@@ 4.5 Ablation Study
To demonstrate the effectiveness of our model and provide experimental and theoretical foundations for future improvements in model design, training methods, data augmentation techniques, and modifications, we conducted an ablation study. This study provided a comprehensive analysis of various factors influencing the performance of our model. The experimental results and theoretical insights obtained from these experiments not only validated the effectiveness of our model but also laid the groundwork for designing improved models, selecting appropriate training methods, devising better data augmentation techniques, and making informed model modifications.
@@ 4.5.1 Different Dataset
We evaluated our model in Schoffelen dataset (Schoffelen et al., 2019) to illustrate our effectiveness. As shown in Tab. 4, our model has comparable performance on two datasets. This illustrates that our model is applicable under many scenarios, including English and Dutch, and can be adapted on different layouts.
@@
4.5.2 Splits
We conducted further experiments using the GWilliams (Gwilliams et al., 2023) dataset to evaluate NeuSpeech’s generalizability. To test the model’s ability to generalize across recording sessions, subjects, and unique sentences, we developed three additional training/validation splits beyond the standard 8:1 split. Specifically, we withheld: 1) all data from the second recording session, 2) all instances of the story “lw1”, and 3) a random 10% sample of unique sentences. For the latter two splits, we ensured no overlap sentences between the training and test sets for the held-out data.
Table 5: Dataset splits.Noise means input model with Gaussian noise. Random means randomly choosing sentences from training set as predictions for testing set. Retrieve means find best match in training set for each test sentence as predictions using sentence similarity using “all-MiniLM-L6-v2” model. Generation means our model’s generated text. Here, we only shows the BLEU-1 score. Note, for hold-out story and sentences, we used nucleus sampling p=0.25 for generation.
Hold-out Noise Random Retrieve Generation
Session 0.0 4.94 100 53.16
Story 0.0 4.97 11.01 6.91
Sentences 0.0 4.79 16.38 6.54
Table 6: Transcription results on different splits, where bold denotes a correct match between ground truth and our prediction. Underline denotes a fuzzy match with similar semantic meanings.
Hold-out sentences
Ground Truth: I looked down and saw he was pointing to another marker just slightly different from the one we had just left
Prediction: It was different from now and it just wasn’t working well.
Hold-out story
Ground Truth: Mahogany black curly hair cut short in the typical military style framed a dark skinned face
Prediction: I must admit, this is a great black or cold weather. It had things such as
As shown in Tab. 5, the model maintained good performance when trained and tested on separate recording sessions, demonstrating its ability to generalize across sessions. NeuSpeech significantly outperformed random baselines on splits with no training-test overlap, validating its generalizability despite low text similarity between train and test sentences.Tab. 6 presents sample outputs, revealing the model can still capture some semantic meaning and words when generalizing, though it struggles to produce fully correct novel sentences.
@@ 4.5.3 Pretrain
To investigate the effectiveness of pretraining, we conducted experiments involving pretraining on one dataset and fine-tuning on the other. Specifically, we pretrained the model on the Schoffelen dataset (Schoffelen et al., 2019) and fine-tuned on GWilliams (Gwilliams et al., 2023), and vice versa. As shown in Table 7, pretraining provided benefits in some settings.
Table 7: The performance of pretraining. The column Pretrain Dataset means trained on GWilliams, Schoffelen and none. Bold for highest socres.
@@
In contrast, fine-tuning the model pretrained on Schoffelen Still allowed for improvements to be made on the larger GWilliams distributions. This suggests pretraining can help in low-data regimes but may inhibit learning when moving to a dataset with a broader sample scope. Overall, these pretraining experiments provide insights into how architectural inductive biases interact with dataset characteristics to influence translation performance.
@@ 4.5.4 Joint Training
The motivation behind conducting this experiment was to explore whether the simplest approach of padding channels could enable the model to learn shared features from both datasets which is multi-language, multi layout and multi subject. It is important to note that the GWilliams dataset comprises only 208 channels, thus we applied padding on the channel dimensions to match the Schoffelen’s channel dimension of 273. We are concerned that this approach might introduce significant interference, rendering both datasets untrainable. However, our experiment revealed that our worries were unfounded. Despite the decrease in model performance caused by the simple channel padding method, it is indeed effective in integrating the two datasets.
@@ 4.5.5 Scaling Law
This finding suggests that scaling up the model size can be a valuable strategy for enhancing the capabilities and effectiveness of the model in this particular task. However, it is important to consider that larger models may suffer from degradation due to the limited availability of data. Therefore, it is crucial to carefully select an appropriate model size.
Additionally, we observe that the number of epochs required to achieve convergence decreases by approximately half as the model size increases. This observation implies that, even under GPU constraints, increasing the size of the model significantly improves performance without imposing a prohibitive training time penalty. Larger models are able to reach convergence more quickly, despite their increased capacity, resulting in training durations comparable to smaller models. By carefully balancing the bias-variance tradeoff for each specific task and dataset, researchers can fine-tune the model to generalize optimally within the limitations of the available data, thereby maximizing decoding quality given the computational resources at hand.
@@ 4.5.6 Data augmentation
There are a lot of research works that used data augmentation in EEG/MEG signal, Bai et al. (Bai et al., 2023) used time-wise masking. Now, we explore three kinds of augmentation, which are masking, noise injecting, and shifting, and compare them with baseline using Whisper-Base in Sec. 4.5.5 which does not have data augmentation. We set three types of masking, which are time-wise, channel-wise, and block-wise. The mask unit is [channel,40],[1,time],[1,40], for the three masking methods respectively. We set two signal-to-noise ratios (SNR), 0dB and 15dB, of Gaussian noise with a probability of 0.5 and 1, which compose 4 settings. We also set 0.5 and 1 probability of data shifting to see if moving the neural data uniformly in the range of max input duration, along with the timestamps can help the model reduce over-fitting. Fig. 2 shows that block masking can improve performance a lot when the mask ratio is relatively low. However, when the masking ratio is bigger, the performance seems to decrease quickly. It can also be noticed that time masking is much worse than the other two mask types to our surprise. It may indicate that the redundancy of the MEG signal is much lower at this sampling rate compared to images which usually use a mask ratio of 75% (He et al., 2022), when the mask ratio is too big to lose essential information, it can be hard for the model to recover information back correctly. Shifting gains no improvements under our experimental setting, oppositely it degrades model performance significantly. To our surprise, adding strong noise which is 0dB can enhance the model’s performance greatly. But for weak noise of SNR 15dB, there is no benefit in our experimental setting. It reminds us that the MEG signal is very noisy, thus using strong noise may let the model better eliminate noise. Therefore, we recommend utilizing block masking with a small mask ratio and low signal-to-noise ratio (SNR) noise for data augmentation as they have demonstrated their effectiveness in enhancing the performance of the models..
@@
Figure 2: Performance changes with different data augmentations. The probability means the likelihood of adding augmentation on each data segment.
4.5.7 Modification
In line with (Défossez et al., 2023), we modified the whisper encoder by replacing the first convolution layer with two separate convolution layers to increase dimensionality and reduce temporal width. Comparing our base model with and without the point-wise convolution layer (Table 9), we observed a significant drop in performance without the layer. This highlights the crucial role of the point-wise convolution layer in achieving good results by capturing important data features and representations.

### v3

@@ 3.4 Generated Samples
Decoding Results on Schoffelen [12]
@@
Table 3: Transcription results comparison with different methods. Bold font means exact match of words. Underline for similar but not exact matching. w/ means with. tf means teacher-forcing, pt means pretraining, jt means joint-training. eeg-to-text [3] here is using beam search as the same setting of ours which is described in Sec. 3.2. More samples are provided in supp. E.
(1) GT I seen him since high school maybe twenty years before and we were never buddies in the first place
NeuSpeech I seen him since high school when I was young, at least before and we were never buddies in any place.
NeuSpeech w/ tf i seen him since high school when eventwenty years before and we were never buddies in onethe first place
NeuSpeech w/ pt I seen him since high school maybe twenty years before and we were never buddies in the first place he talked to me since t.
NeuSpeech w/ jt I seen him since high school maybe twenty years before and we were never buddies in the first place
eeg-to-text [3] It was a long time since I had last seen him in the flesh
eeg-to-text [3] w/ tf was the and I school a years ago that he had both close the first place
@@
eeg-to-text [3] He said he had no idea how long it would take him to get back home
eeg-to-text [3] w/ tf father was running gone I was ready to the dark drive up Iacia me the window said me to was a a he was looking for
@@
eeg-to-text [3] It was the first time I had seen him in a long time and it made me very nervous
eeg-to-text [3] w/ tf little then he the seat and the place go it take it he one man would you
In Tab. 2, we present visual examples of text that have been generated from MEG signals. The training of Schoffelen [12] is described in Sec. 3.5.1 These examples showcase the capabilities of our model in generating meaningful text, despite the difficulties associated with translating thoughts and the limited prior research in this area. While our model may not yet achieve the same level of performance as nowadays speech recognition tasks, it demonstrates the ability to align keywords and construct sentences with similar structures. This is a promising indication of the potential of our approach in the field of brain-to-text translation.
@@
We have provided Tab. 3 below, which compares the results of different approaches, namely NeuSpeech, NeuSpeech with teacher forcing, NeuSpeech with pre-training, and joint training. We observe that using NeuSpeech alone often result in semantically coherent sentences, even if not all the words were correct. Particularly in example (3), there was even a hint of expressive tone. However, when NeuSpeech was generating text with teacher forcing, although there were more correct word matches, there were often concatenated words, such as "onethe." The reason for this occurrence is not yet clear, but it significantly affects the semantic coherence of the sentences.
@@
In contrast, the joint training model seems to tolerate longer sequences of word errors. For example, in (3), the model does not predict correct words for most of the sentence but accurately predicts the last eleven words, which is neglected by the higher-scoring pre-trained model. This indicates that, even for a model that reaches a certain performance level, we cannot solely rely on scoring metrics. Each model may have unique attributes that we can exploit. In addition, previous approaches for EEG-to-text translation seem to be less effective, even with the use of teacher forcing, they produced only a few words that match the true labels. This further highlights the superiority of our approach. In addition, the EEG-to-text model always predicts concepts like a long time, which is seen in supp.listing 4.
In summary, the analysis of generated samples from different approaches reveals valuable insights, which emphasize the importance of considering both quantitative metrics and understanding the unique properties and strengths of different models. Furthermore, the comparison with previous EEG-to-text [3] methods underscores the advancements achieved by our approach.
@@ 3.5 Ablation Study
To demonstrate the effectiveness of our model and provide insight for large corpus training on this task. This study provided a comprehensive analysis of various factors influencing the performance of our model. The experimental results and theoretical insights obtained from these experiments not only validated the effectiveness of our model but also laid the groundwork for designing improved models, selecting appropriate training methods, devising better data augmentation techniques, and making informed model modifications.
@@ 3.5.1 Different Dataset
We evaluated our model in Schoffelen dataset [12] to illustrate our effectiveness. As shown in Tab. 4, our model has comparable performance on two datasets. This illustrates that our model is applicable under many scenarios, including English and Dutch, and can be adapted on different layouts.
@@ 3.5.2 Pretrain
Table 5: The performance of pretraining. The column Pretrain Dataset means trained on GWilliams, Schoffelen, and none. Bold for highest socres.
@@
Here pretrain means training on one MEG dataset, and continuing training on another MEG dataset. Specifically, we trained the model on the Schoffelen dataset [12] and fine-tuned on GWilliams [11], and vice versa. As shown in Table 5, pretraining provided benefits in some settings.
@@
In contrast, fine-tuning the model pre-trained on Schoffelen Still allowed for improvements to be made on the larger GWilliams distributions. This suggests pretraining can help in low-data regimes but may inhibit learning when moving to a dataset with a broader sample scope. Overall, these pretraining experiments provide insights into how architectural inductive biases interact with dataset characteristics to influence translation performance.
@@ 3.5.3 Joint Training
Joint training means training with multiple datasets at one time. We want to find out if the model can learn from multiple datasets together which is a key to enable large-pretraining of models. It is important to note that the GWilliams dataset comprises only 208 channels, thus we used the simplest method of applying padding on the channel dimensions to match the Schoffelen’s channel dimension of 273 to see if this simple method works out.
@@ 3.5.4 Scaling Law
Additionally, we observe that the number of epochs required to achieve convergence decreases by approximately half as the model size increases. By carefully balancing the bias-variance trade-off for each specific task and dataset, researchers can fine-tune the model to generalize optimally within the limitations of the available data, thereby maximizing decoding quality given the computational resources at hand.
@@ 3.5.5 Data augmentation
There are a lot of research works that used data augmentation in EEG/MEG signal, Bai et al. [17] used time-wise masking. Now, we explore three kinds of augmentation, which are masking, noise injecting, and shifting, and compare them with baseline using Whisper-Base in Sec. 3.5.4 which does not have data augmentation. We set three types of masking, which are time-wise, channel-wise, and block-wise. The mask unit is [channel,40],[1, time],[1,40], for the three masking methods respectively. We set two signal-to-noise ratios (SNR), 0dB and 15dB, of Gaussian noise with a probability of 0.5 and 1, which compose 4 settings. We also set 0.5 and 1 probability of data shifting to see if moving the neural data uniformly in the range of max input duration, along with the timestamps can help the model reduce over-fitting.
Fig. 3 shows that block masking can improve performance a lot when the mask ratio is relatively low. However, when the masking ratio is bigger, the performance seems to decrease quickly.
@@
Figure 3: Performance changes with different data augmentations. The probability means the likelihood of adding augmentation to each data segment.
It can also be noticed that time masking is much worse than the other two mask types to our surprise. It may indicate that the redundancy of the MEG signal is much lower at this sampling rate compared to images which usually use a mask ratio of 75% [18], when the mask ratio is too big to lose essential information, it can be hard for the model to recover information back correctly. Shifting gains no improvements under our experimental setting, oppositely it degrades model performance significantly. To our surprise, adding strong noise which is 0dB can enhance the model’s performance greatly. But for weak noise of SNR 15dB, there is no benefit in our experimental setting. It reminds us that the MEG signal is very noisy, thus using strong noise may let the model better eliminate noise. Therefore, we recommend utilizing block masking with a small mask ratio and low signal-to-noise ratio (SNR) noise for data augmentation as they have demonstrated their effectiveness in enhancing the performance of the models.
3.5.6 Data ratio and Fine-tune layers
Figure 4: Fine-tune layers and data ratio. Blue line for different fine-tune layers, green line for different training data ratios. Note the whisper-base model only has 6 layers.
We explore the ratio and data used in training, locking validation set and test set to ensure fairness. In Fig. 4, we can see that performance rises as data is added as expected. Higher performance may be achieved if more data is provided since the curve seems far behind reaching the plateau.
It can be observed that the more layers we fine-tune, the better the model will be. Note, that it is only on this base model, and may not be applicable to all.
4 Limitations
While NeuSpeech shows promising results on benchmark MEG-to-text tasks, there remain opportunities to strengthen the scientific validity and real-world applicability of the framework. Future work could focus on incorporating less structured brain signals to better validate the translatability of direct human thought, including physical sensor position data for improved accuracy, and exploring augmented datasets to address data scarcity issues that currently limit the model’s ability to generalize and generate completely correct novel sentences. Addressing these limitations through collecting additional training data and enriching the inputs could help advance this emerging field by propelling NeuSpeech towards decoding thought from brain signals in varied real-world scenarios.
5 Conclusion
This paper explores MEG-to-text using existing speech recognition training, introducing the concept of decoding neural signals as speech. This approach brings high performance on two datasets on standard metrics and stretches to decode MEG signals without using teacher forcing. We also demonstrate the feasibility of pre-training or jointly training large neural speech models across layouts and languages. We have also explored the effects of augmentation and so on. Despite these advances, the results of hold-out scenarios are relatively low due to the limited number of sentences and the noisy nature of the MEG signal. In our future work, we will solve the problem of limited training sets and also find more mechanisms to fuse multi-modalities, and multi-layers with physical information embeddings.
Supplementary Material for
NeuSpeech: Decode Neural signal as Speech
Appendix A Dataset
The Gwilliams [11] dataset is created by Laura Gwilliams and colleagues. Below is the dataset detail quoted from their paper.
Participants Twenty-seven English-speaking adults were recruited from the subject pool of NYU Abu Dhabi (15 females; age: M=24.8, SD=6.4). All participants provided written informed consent and were compensated for their time. Participants reported having normal hearing and no history of neurological disorders. All but one participants (S20) were native English speakers. All but five participants (S3, S11, S15, S19, S20) performed two identical one-hour-long sessions. These two recording sessions were separated by at least one day and at most two months depending on the availability of the experimenters and of the participants. The study was approved by the Institution Review Board (IRB) ethics committee of New York University Abu Dhabi.
Procedure Within each  1 h recording session, participants were recorded with a 208 axial-gradiometer MEG scanner built by the Kanazawa Institute of Technology (KIT), and sampled at 1,000 Hz, and online band-pass filtered between 0.01 and 200 Hz while they listened to four distinct stories through binaural tube earphones (Aero Technologies), at a mean level of 70 dB sound pressure level. Before the experiment, participants were exposed to 20 sec of each of the distinct speaker voices used in the study to (i) clarify the structure of the session and (ii) familiarize the participants with these voices. The order in which the four stories were presented was assigned pseudo-randomly, thanks to a "Latin-square design" across participants. This participant-specific order was used for both recording sessions. To ensure that the participants were attentive to the stories, they answered, every  3 min and with a button press, a two- alternative forced-choice question relative to the story content (e.g. ‘What precious material had Chuck found? Diamonds or Gold’). Participants performed this task with an average accuracy of 98%, confirming their engagement with and comprehension of the stories. Participants who did not already have a T1-weighted anatomical scan usable for the present study were scanned in a 3T Magnetic-Resonance-Imaging (MRI) scanner after the MEG recording to avoid magnetic artefacts. Six participants did not return for their T1 scan. Before each MEG session, the head shape of each participant was digitized with a hand-held FastSCAN laser scanner (Polhemus), and co-registered with five head-position coils. The positions of these coils with regard to the MEG sensors were collected before and after each recording and stored in the ’marker’ file, following the KIT’s system. The experimenter continuously monitored head position during the acquisition to ensure that the participants did not move.
Stimuli Four English fictional stories were selected from the Open American National Corpus – a manually annotated corpus distributed without license or other restrictions1:
• ‘Cable spool boy‘: a 1,948-word story narrating two young brothers playing in the woods
• ‘LW1‘: a 861-word story narrating an alien spaceship trying to find its way home
• ‘Black willow‘: a 4,652-word story narrating the difficulties an author encounters during writing.
• ‘Easy money‘: a 3,541-word fiction narrating two friends using a magical trick to make money.
An audio track corresponding to each of these stories was synthesized using Mac OS Mojave’s (c) text-to-speech. To help decorrelate language from acoustic representations, we varied both voices and speech rate every 5-20 sentences. Specifically, we used three distinct synthetic voices, namely ‘Ava‘, ‘Samantha‘ and ‘Allison‘ speaking between 145 and 205 words per minute. Additionally, we varied the silence between sentences between 0 and 1,000 ms. Each story was divided into around 5 min sound files. In between these sounds – approximately every 30 s – we played a random word list generated from the unique content words (nouns, proper nouns, verbs, adverbs and adjectives) of the preceding 5 min segment presented in random order. In addition, a very small fraction (<1%) of non-words were introduced in natural sentences. Hereafter, and following the BIDS labeling, each "task" corresponds to the concatenation of these sentences and word lists. Each subject listened to the exact same set of four tasks, only in a different block order.
The Schoffelen dataset is a multimodal neuroimaging dataset, below is the detailed description from their paper [12]. Note we only used auditory sentence in our experiments.
Subjects A total of 204 native speakers of Dutch (100 males) with a mean age of 22 years (range: 18 to 33 years) were included in the study. In the informed consent procedure, they explicitly consented for the anonymized collected data to be used for research purposes by other researchers. The subjects took part in both the fMRI and MEG part of the study, in a counterbalanced fashion. Each subject performed the task in either the visual or the auditory modality. All subjects were right-handed, had normal or corrected-to-normal vision, and reported no history of neurological, developmental or language deficits. The study was approved by the local ethics committee (CMO – the local “Committee on Research Involving Human Subjects” in the Arnhem-Nijmegen region) and followed guidelines of the Helsinki declaration.
Stimulus material The total stimulus set consisted of 360 sentences in Dutch, and their scrambled word list counterparts. The sentences consisted of two types: 180 of the sentences contained a relative clause (RC+), to create a more difficult syntactic structure. The other 180 sentences consisted of a main clause and a simple subordinate clause (RC-), to create an easier structure. The word lists were created by scrambling the words from the sentences such that three or more consecutive words did not form a coherent fragment. For an example of the sentences see Table 1. All sentences varied between 9 and 15 words in length.
Each subject was presented with a subset of 180 sentence stimuli, and 180 word list stimuli, where we ensured that, for a given sentence, they were not exposed to the corresponding word list. During the task MEG part of the experiment, the subjects were presented with 120 sentences and 120 word lists, and during the fMRI part of the experiment they were presented with 60 sentences and 60 word lists. Across subjects, all stimuli were presented the same number of times in the sentence and in the word list condition.
Each sentence and corresponding word list contained a noun that was at the same ordinal position, which varied between the third and thirteenth word position (which is denoted here as the ‘target word’). This allows for psychometrically controlled comparisons between the conditions using these isolated nouns. These words were matched in terms of lexical frequency and word length. To control for similarity in context, the word preceding the target did not differ in word length by more than two letters between a sentence and its word list counterpart. The word frequency was based on the lemma frequency according to the SUBTLEX-NL database of Dutch word frequencies21, and was on average 27,6/million (SD = 62,1/million). All stimulus material is shared along with the neuroimaging data.
The auditory versions of the stimuli were recorded by a female native Dutch speaker. The word lists were pronounced with a neutral prosody and a clear pause in between each word. The audio files were recorded in stereo at 44100 Hz. During the post processing the audio files were low-pass filtered at 8500 Hz and normalized such that all audio files had the same peak amplitude, and same peak intensity. In the word list condition, each word was separated by 300 ms of total silence. The transition from silence to speech was ramped at the onset (rise time of 10 ms) and offset (fall time of 10 ms) of single words in the word list condition, and for sentence onset. The onset of the first and target word vocalizations were determined manually for each audio file, using the Praat software (http://www.praat.org, RRID:SCR_016564).
Experimental Design and Procedure The total stimulus set was divided into two sets of 180, and each of these were subsequently divided into three subsets. Each subject was presented with 2/3 of the stimuli set in the MEG (120 trials of each condition) and 1/3 in the fMRI (60 trials). Across subjects, each subset was presented as many times in MEG as in fMRI. Different subjects that were presented with the same subset had the stimuli presented in a different (randomized) order. Within an experimental session, the stimuli were presented in blocks, alternating between sentence blocks (containing 5 sentences) and word list blocks (containing 5 word lists), for a total of 24 (MEG) or 12 (fMRI) blocks. The starting block type (either sentences or word list) was randomized across subjects.
In order to check for compliance, 20% of the trials were followed by a ‘Yes’/‘No’ question about the content of the previous sentence/word list. Half of the questions on the sentences addressed the content of the sentence (e.g. Did grandma give a cookie to the girl?) whereas the other half, and all of the questions about the word lists, addressed one of the main content words (e.g. Was the word ‘grandma’ mentioned?). A substantial part of the questions on complex relative clause (RC+) sentences concerned content from the relative clause. Subjects answered the question by pressing a button for ‘Yes’/‘No’ with their left index and middle finger, respectively.
At the start of each block there was a 1500 ms presentation of the block type: zinnen (sentences) or woorden (words). In sentences, the first word began with a capital letter, and the last word end with a full stop. The inter-trial interval was jittered between 3200–4200 ms. During this period, an empty screen was presented, followed by a fixation cross.
Stimuli were presented using the Presentation software (Version 16.0, Neurobehavioral Systems, Inc).
Prior to the task, subjects read a written instruction of the task and were allowed to ask questions for clarification. Furthermore, the experimenter emphasized that the sentences and word lists should be attended carefully, and discouraged attempts to integrate the words in the word list condition. Finally, to familiarize the subjects with the task, they did a practice task with stimuli separate from the actual study task. Figure 1 shows a schematic overview of the study procedure.
Auditory language task In the auditory task the stimuli were presented via plastic tubes and ear pieces to both ears. Before the experiment, the hearing threshold was determined individually and the stimuli were then presented at an intensity of 50 dB above the hearing threshold. In fMRI, the hearing threshold was determined on top of the EPI-sequence noise, to verify that all stimuli were clearly audible.
We arranged our preprocessed data into .jsonl files, here is an example.
1{"speech": {"path": "/hpc2hdd/home/yyang937/datasets/gwilliams2023/preprocess5/sub-21/ses-0/meg/sub-21_ses-0_task-3_senid_11_audio.wav", "sr": 16000}, "eeg": {"path": "/hpc2hdd/home/yyang937/datasets/gwilliams2023/preprocess5/sub-21/ses-0/meg/sub-21_ses-0_task-3_senid_11_meg.npy", "sr": 200}, "duration": 3.700000000000003, "language": "English", "sentence": "Although the concept was a simple one Allan thought it had potential", "sentences": [{"text": "Although the concept was a simple one Allan thought it had potential", "start": 0.0, "end": 3.700000000000003, "duration": 3.700000000000003, "words": [{"word": "Although", "start": 24.653999999999996, "end": 24.964}, {"word": "the", "start": 24.964, "end": 25.073999999999998}, {"word": "concept", "start": 25.073999999999998, "end": 25.593999999999994}, {"word": "was", "start": 25.593999999999994, "end": 25.733999999999995}, {"word": "a", "start": 25.733999999999995, "end": 25.803999999999988}, {"word": "simple", "start": 25.804000000000002, "end": 26.12400000000001}, {"word": "one", "start": 26.423999999999992, "end": 26.774}, {"word": "Allan", "start": 26.933999999999997, "end": 27.233999999999995}, {"word": "thought", "start": 27.233999999999995, "end": 27.453999999999994}, {"word": "it", "start": 27.453999999999994, "end": 27.533999999999992}, {"word": "had", "start": 27.543999999999997, "end": 27.72399999999999}, {"word": "potential", "start": 27.724000000000004, "end": 28.354000000000013}]}], "subj": 21, "story": "The_Black_Willow", "story_id": 3.0, "seq_id": 11.0, "sound_id": 0.0, "speech_rate": 205.0, "voice": "Samantha", "start": 92.824, "end": 96.52400000000002, "audio_start": 68.17, "audio_end": 71.87}
Appendix B Splits
Table 7: Different dataset splits. Noise means input model with Gaussian noise with same length of MEG signal. We shows the BLEU-1 score and an example of generation. Underline denotes a fuzzy match with similar semantic meanings. Note, for hold-out story and sentences, we used nucleus sampling p=0.25 for generation.
Hold-out Noise Generation and example sentence
Session 0.0 53.16 Ground truth: something worth reading worth writing
Prediction: some kind of what i sell
Story 0.0 6.91 Ground truth: I looked down and saw he was pointing to another marker just slightly different from the one we had just left
Prediction: It was different from now and it just wasn’t working well.
Sentences 0.0 6.54 Ground truth: her first officer was staring at his panel
Prediction: he glared at allan eyes him and demanding a response but he
We conducted further experiments using the GWilliams [11] dataset to evaluate NeuSpeech’s generalizability. To test the model’s ability to generalize across recording sessions, subjects, and unique sentences, we developed three additional training/validation splits beyond the standard 8:1 split. Specifically, we withheld: 1) all data from the second recording session, 2) all instances of the story “lw1", and 3) a random 10% sample of unique sentences. For the latter two splits, we ensured no overlap sentences between the training and test sets for the held-out data.
As shown in Tab. 7, the model maintained good performance when trained and tested on separate recording sessions, demonstrating its ability to generalize across sessions. NeuSpeech significantly outperformed random baselines on splits with no training-test overlap, validating its generalizability despite low text similarity between train and test sentences. Also we present sample outputs, revealing the model can still capture some semantic meaning and words when generalizing, though it struggles to produce fully correct novel sentences.
Appendix C Evaluation on words
To figure out if our model is learning MEG signal, we designed this experiment. We randomly clip segments of varying lengths and use the corresponding words to train, and then we test on every single words. It turns out our model gets 5.8% accuracy, since the test set has 2251 words, the chance level is 0.04%, our model’s word classification accuracy is 130 times the chance level. This means our model has learned something.
Appendix D Modification